20-[4+8y-4(y+3)]=-3(5y-6)-[9(y-1)-14y+11]
i keep on getting -8

Answer by jim_thompson5910(35256)

20-[4+8y-4(y+3)]=-3(5y-6)-[9(y-1)-14y+11]

20-[4+8y-4y-12]=-3(5y-6)-[9y-9-14y+11]

20-[4y-8]=-3(5y-6)-[-5y+2]

20-4y+8=-15y+18+5y-2

-4y+28=-10y+16

-4y+28+10y=-10y+16+10y

-4y+28+10y=16

-4y+28+10y-28=16-28

-4y+10y=16-28

6y=-12

6y/6=-12/6

y = -2

So the solution is y = -2

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------